P(E)=1-P(E^C)4PermutationHeartContinuousClubDiamondCombinationMultiplyDiscreteprobabilityproportion,percentnormalcdf13binomialpdfbinomialcdf26Atleast12invNormAddSpadeOrAndP(E)=1-P(E^C)4PermutationHeartContinuousClubDiamondCombinationMultiplyDiscreteprobabilityproportion,percentnormalcdf13binomialpdfbinomialcdf26Atleast12invNormAddSpadeOrAnd

Post Exam 2 - Call List

(Print) Use this randomly generated list as your call list when playing the game. There is no need to say the BINGO column name. Place some kind of mark (like an X, a checkmark, a dot, tally mark, etc) on each cell as you announce it, to keep track. You can also cut out each item, place them in a bag and pull words from the bag.


1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
  1. P(E)=1-P(E^C)
  2. 4
  3. Permutation
  4. Heart
  5. Continuous
  6. Club
  7. Diamond
  8. Combination
  9. Multiply
  10. Discrete
  11. probability proportion,percent
  12. normalcdf
  13. 13
  14. binomialpdf
  15. binomialcdf
  16. 26
  17. At least
  18. 12
  19. invNorm
  20. Add
  21. Spade
  22. Or
  23. And